OpenStudy (anonymous):

Evaluate the integral ∫ (7dx)/(√16-49x^2) Is it ok if I simplify it to ∫(7dx)/(4-7x) ??

OpenStudy (anonymous):

No it cannot be simplified that way

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 7 }{ \sqrt{16-49x^2} }dx\]

OpenStudy (anonymous):

T.T ok

OpenStudy (anonymous):

By doing that I don't know if the square root would go away though or would it?? :/

OpenStudy (anonymous):

Ya the square root disappears\[(x^a)^b=x^{ab}\]\[\large (x^{\frac{ 1 }{ 2 }})^2=x^1=x\]

OpenStudy (anonymous):

I don't think I can do that, hmmm...

OpenStudy (anonymous):

oh ok :/ I'll try to google this...lets see

jimthompson5910 (jim_thompson5910):

\[\Large \frac{7}{\sqrt{16 - 49x^2}}\] \[\Large \frac{7}{\sqrt{16(1 - \frac{49}{16}x^2)}}\] \[\Large \frac{7}{\sqrt{16}\sqrt{1 - \frac{49}{16}x^2}}\] \[\Large \frac{7}{4\sqrt{1 - \frac{49}{16}x^2}}\] \[\Large \frac{7}{4\sqrt{1 - \left(\frac{7x}{4}\right)^2}}\]

jimthompson5910 (jim_thompson5910):

so \[\Large \int\frac{7dx}{\sqrt{16 - 49x^2}}\] is the same as \[\Large \int\frac{7dx}{4\sqrt{1 - \left(\frac{7x}{4}\right)^2}}\] \[\Large \frac{7}{4}\int\frac{dx}{\sqrt{1 - \left(\frac{7x}{4}\right)^2}}\] \[\Large \frac{7}{4}\int\frac{4du}{7\sqrt{1 - u^2}}\] \[\Large \frac{7}{4}*\frac{4}{7}\int\frac{du}{\sqrt{1 - u^2}}\] \[\Large \int\frac{du}{\sqrt{1 - u^2}}\] where u = 7x/4 ---> du = 7dx/4 ---> dx = 4du/7

OpenStudy (anonymous):

Wow! Thank you so much! Both of you thanks! =)

jimthompson5910 (jim_thompson5910):

from there, you use the idea that if y = arcsin(x), then dy/dx = 1/sqrt(1 - x^2)

jimthompson5910 (jim_thompson5910):

but you do it in reverse

jimthompson5910 (jim_thompson5910):

once you've integrated, don't forget to replace u with what it is in terms of x

jimthompson5910 (jim_thompson5910):

oh and i'm sure you know this, but don't forget the +C (it gets me everytime)

OpenStudy (anonymous):

yes thank you so much! I was doing it out right now, I'll post the answer once I'm done. :)

jimthompson5910 (jim_thompson5910):

ok, yw

OpenStudy (anonymous):

hmm...I got[ (1-u^2)^1.5]/1.5 + c but I think its wrong did I have to include the du=7dx/4?? It's probably all wrong...

OpenStudy (anonymous):

QUESTION!! Is this correct? \[7\int\limits \frac{ dx }{ \sqrt{49}\times \sqrt{\frac{ 16 }{ 49}-x^2}} \] \[= \frac{ 7 }{ 7 }\int\limits \frac{ dx }{ \sqrt{\frac{ 9 }{ 7 }^2 -x^2} }\]

OpenStudy (anonymous):

*BTW the numbers that look like 0 are 9

jimthompson5910 (jim_thompson5910):

no you should have arcsine in your answer

jimthompson5910 (jim_thompson5910):

\[\Large \int\frac{du}{\sqrt{1 - u^2}} = \arcsin(u) + C\]

jimthompson5910 (jim_thompson5910):

\[\Large \int\frac{du}{\sqrt{1 - u^2}} = \arcsin(u) + C\] \[\Large \int\frac{du}{\sqrt{1 - \left(\frac{7x}{4}\right)^2}} = \arcsin\left(\frac{7x}{4}\right) + C\] \[\Large \int\frac{7dx}{\sqrt{16 - 49x^2}} = \arcsin\left(\frac{7x}{4}\right) + C\]

OpenStudy (anonymous):

would the answer be: \[\sin ^{-1} (\frac{ x }{ \frac{ 9 }{ 7 } }) +c\]

jimthompson5910 (jim_thompson5910):

i agree with the sine inverse, but not sure how you got the 9/7

OpenStudy (anonymous):

oops would the final answer be: \[\sin ^{-1} (7/4 x)+c\]

jimthompson5910 (jim_thompson5910):

yep, that's what I wrote above arcsin = sin^(-1)

OpenStudy (anonymous):

ohh lolol ^_^ ok TY!!

jimthompson5910 (jim_thompson5910):

np